A hydraulic lift consists of two pistons that connect to each other by an incompressible fluid. If one piston has an area of 0.15 m² and the other an area of 6.0 m², how large a mass can be raised by a force of 130 N exerted on the smaller piston?
1
Expert's answer
2020-11-05T10:39:27-0500

The presure exerted on the smaller piston is equal to the preasure provided by the big one:


"\dfrac{F_1}{A_1} = \dfrac{F_2}{A_2}"

where "F_1 = 130N" is the foce exerted on the smaller piston, "A_1 = 0.15m^2" is the area of the smaller piston, "F_2" is the force provided by the big piston, and "A_2 = 6m^2" is the area of the big piston. Expressing "F_2", get:


"F_2 = F_1\dfrac{A_2}{A_1} =130\times \dfrac{6}{0.15} = 5200N"

A mass that can be raised by this force is:


"m = \dfrac{F_2}{g} = \dfrac{5200N}{9.81m/s^2} \approx 530.07kg"

Answer. 530.07 kg.